# 3D ISM structure challenges the Serkowski relation

### v1

@@ Abstract
Context. The Serkowski relation is the cornerstone of studies of starlight polarizations as a function of wavelength. Although empirical, its extensive use since its inception by Serkowski et al. (1975) to describe polarization induced by interstellar dust has elevated the relation to the status of an indisputable “law”: this is the benchmark against which models of interstellar dust grains are validated.
Aims. We revisit the effects of the three - dimensional structure of the interstellar medium (ISM) on the wavelength dependence of interstellar polarization.
Methods. We use analytical models to show how the wavelength dependence of both the polarization fraction and direction is affected by the presence of multiple clouds along the line of sight (LOS). We account for recent developments in dust distribution modelling and we utilize an expanded archive of stellar polarization measurements to establish the effect of multiple clouds along the LOS. We highlight concrete examples of stars whose polarization profiles are severely affected by LOS variations of the dust grain and magnetic field properties, and we provide a recipe to accurately fit multiple cloud Serkowski models to such cases.
Results. We present, for the first time, compelling observational evidence that the three - dimensional structure of the magnetized ISM often results to the violation of the Serkowski relation. We show that three - dimensional effects impact interstellar cloud parameters derived from Serkowski fits. In particular, the dust size distribution in single - cloud sightlines may differ from analyses that ignore 3D effects, with important implications for dust modelling in the Galaxy.
Conclusions. Our results suggest that multiwavelength stellar polarization measurements offer an independent probe of the LOS variations of the magnetic field orientation, and thus constitute a potentially valuable new tool for the 3D cartography of the interstellar medium. Finally, we caution that, unless 3D effects are explicitly accounted for, a poor fit to the Serkowski relation does not, by itself, constitute conclusive evidence that a star is intrinsically polarized.
@@ 1 Introduction
Dust is ubiquitous in the interstellar medium (ISM) of our Galaxy and plays a significant role in many astrophysical processes. The thermal (gray-body) emission of interstellar dust spans many frequencies (100 – 1000 GHz), making dust a powerful tracer of the morphology of ISM structures as well as a diagnostic of their physical properties, such as temperature (Hildebrand, 1983; André et al., 2010). At the same time dust emission poses a challenge for cosmological experiments because it veils the cosmic microwave background signal over a wide range of frequencies (Planck Collaboration et al., 2016). Due to the multifaceted role of interstellar dust in astrophysics and cosmology, a detailed understanding of its nature is imperative.
The interaction of aspherical dust grains with the magnetic field permeating the ISM leads to their alignment, resulting in the polarization of light, which can be observed from ultraviolet to sub-mm wavelengths (Andersson et al., 2015). Short-wavelength (UV, optical, near-IR) polarization is induced by dichroic extinction of the starlight (which usually starts out unpolarized) by aligned dust grains. The same grains emit long-wavelength (far-IR, and sub-mm) polarized thermal radiation. Dust polarization provides one of the most reliable and widely employed methods for tracing the properties of the ISM magnetic fields (e.g., Skalidis & Tassis, 2021; Skalidis et al., 2021, 2022; Pattle et al., 2023; Pelgrims et al., 2023, 2024).
Dust-induced stellar polarization follows a complex empirical relation with wavelength, known as the “Serkowski relation” (Serkowski et al., 1975). It expresses the polarization fraction $P$ as a function of wavelength $\lambda$ through the formula
@@
with the use of three free parameters: $P_{max}$ the maximum polarization fraction observed at $\lambda_{max}$, and $K$ that quantifies the spread of the profile. The polarization angle (also referred to as electric vector position angle, EVPA, denoted here as $\theta$) is implicitly assumed to be constant with wavelength.
@@
Since its discovery by Serkowski et al. (1975), this relation has been extensively studied and confirmed by a number of authors (Wilking et al., 1980; Whittet et al., 1992, 2001; Bagnulo et al., 2017; Cikota et al., 2018). It is so well established that it is often dubbed the “Serkowski law”. It is even a standard practice in the literature to use it to extrapolate the polarization fraction in some wavelength from measurements made at another wavelength (Bijas et al., 2024; Neha et al., 2024; Ginski et al., 2024). Most importantly, the Serkowski relation with its complicated shape provides stringent constraints for dust grain models (Guillet et al., 2018; Hensley & Draine, 2021; Draine, 2024), while in emission, the degree of polarization is relatively flat across frequencies (Hensley & Draine, 2023).
@@
The Serkowski relation applies for single-cloud lines of sight (LOSs), and to date, whenever it is applied, it is implicitly assumed that the observed (integrated) $P(\lambda)$ signal is induced by a single dominant polarization screen (cloud) along the LOS. Reality, however, is more complicated. Gaia has opened up the third dimension of the sky (depth, Gaia Collaboration et al., 2018), and since then it has become clear that in the majority of the LOSs there are more than one clouds with comparable contributions in the total dust column (e.g., Green et al., 2019; Edenhofer et al., 2024). The variations of the magnetic field along the LOS are imprinted on the individual stellar polarization measurements, allowing the tomographic mapping of the ISM magnetic field (Pelgrims et al., 2024).
@@
The impact of multiple polarizing clouds along the LOS on stellar polarization has been discussed in the literature, though not extensively. Most of the focus has been on the effect on the EVPA, with less attention given to the polarization fraction spectral profile. During the early stages of interstellar polarization (ISP) studies, when the notion of a wavelength-independent EVPA emerged from observational evidence, Treanor (1963) briefly mentioned the possibility of a non-constant EVPA in scenarios involving multiple clouds with different magnetic field alignment and/or different grain sizes. Coyne & Gehrels (1966) attributed the observed variability in EVPA with wavelength among their targets to the potential presence of multiple clouds along the LOS. Martin (1974) was the first to construct mathematical models to describe such scenarios, involving multiple clouds with differing properties along the LOS. Some authors focused on the impact of multiple dust components in the LOS on the parameters derived from Eq. 1. For example, Codina-Landaberry & Magalhaes (1976) explored the behavior of the $K$ parameter in a medium with changing grain alignment, while Clarke & Al-Roubaie (1984) investigated how the presence of two clouds along the LOS would influence the parameters $K$ and $\lambda_{max}$ with simplistic models, finding no firm relationship between their models and contemporary observations.
@@
While modern researchers occasionally reference these earlier works most of the focus remains on the effect of a two-cloud scenario on the EVPA (e.g., McMillan & Tapia, 1977; Messinger et al., 1997; Clarke, 2010; Whittet, 2015; Bagnulo et al., 2017; Whittet, 2022). Little attention is given to the broader consequences of such scenarios, in the $P-\lambda$ profile [^1] Patat et al. (2010) briefly hints on this in Appendix B or the derived dust parameters.
@@
In this article, we revisit the effects of the three-dimensional structure of the ISM on the wavelength dependence of ISP, accounting for recent developments in dust distribution modelling and utilizing an expanded archive of polarimetric measurements. We use analytical models to show how both the $P-\lambda$ and $\theta-\lambda$ profiles are expected to behave in scenarios with multiple clouds along the LOS. For the first time, we present compelling observational evidence to demonstrate these effects and disentangle the polarizing components within the LOS. Ultimately, we demonstrate that the LOS variations of dust and magnetic field properties affect the Serkowski relation in two ways: 1) they challenge the universality of the relation, because Eq. 1 can be a poor fit to the data when 3D effects are significant; and 2) even when Eq. 1 fits the data well, the best-fit parameters may not be representative of dust physics. Conversely, deviations of stellar polarization spectral profiles could potentially be utilized as a diagnostic of the 3D complexity of the magnetized ISM.
@@ 2 Theoretical expectations
In the case of multiple dust clouds along the LOS, initially unpolarized star light is affected sequentially by all of them, until it reaches the observer. In the limit of low polarization the normalized Stokes parameters $q=Q/I$, $u=U/I$ of the transversing light, are additive (e.g. Patat et al., 2010). Therefore the observed $q,u$ in a given wavelength of an intrinsically unpolarized source lying behind $N_{C}$ number of clouds will be
$q_{obs}=\sum^{N_{C}}_{i=1}q_{i}$ (2)
and similarily for $u$, with $q_{i}$ being the $q$ induced by the $i-th$ dust cloud in the LOS.
In order to explore this effect, we created analytical models for the cases of 1 and 2 dust clouds along the LOS, for different parameters of the clouds. Given that the Serkowski relation is an intrinsic property of individual ISM clouds, variations in grain sizes would be manifested by variations in $\lambda_{max}$, which is proportional to the average size of aligned grains (Draine, 2024). Variations in the plane-of-sky (POS) morphology of the magnetic field are traceable through the polarization angle $\theta$. Variations in $\lambda_{max}$ and in $\theta$ between two or more clouds along the LOS significantly affect the cumulative (observed) polarization signal. For every cloud $i$ in every model of Fig. 1, we used Eq.1 to calculated $P_{i}(\lambda)$ for $0.1\mu m\leq\lambda\leq 1.2\mu m$, selecting typical values for $P_{max}$ and $\lambda_{max}$ and fixing $K=1.68\cdot\lambda_{max}$ (Wilking et al., 1980). We also assigned a $\theta_{i}$ value (constant with wavelength) representing the average POS magnetic field direction within the cloud. Thus, for each cloud $i$ we have $q_{i}(\lambda)=P_{i}(\lambda)\cdot\cos{2\theta_{i}}$ and $u_{i}(\lambda)=P_{i}(\lambda)\cdot\sin{2\theta_{i}}$. For the combined model, we calculated $q_{obs}(\lambda)$, $u_{obs}(\lambda)$ from Eq. 2 and finally $\hat{P}(\lambda)$, $\hat{\theta}(\lambda)$ as
@@
$\hat{\theta}(\lambda)=\frac{1}{2}\cdot\tan^{-1}\left(\frac{u_{obs}(\lambda)}{q_{obs}(\lambda)}\right).$ (3b)
@@
The simple case of a single dust cloud (Serkowski relation - Eq.1) is shown in the left panel of Fig. 1 for reference. Two cases with 2 clouds in the LOS are shown in the middle and right panels of Fig. 1. Specifically, the middle panel demonstrates an example where the $P-\lambda$ curve resembles the Serkowski relation, yet the $\theta-\lambda$ profile deviates from constancy. The right panel demonstrates an example where both the $P-\lambda$ and the $\theta-\lambda$ profiles deviate from the Serkowski curves.
@@
Assuming a difference of a factor of two in $\lambda_{max}$ between our two hypothetical clouds yields a bimodal $P(\lambda)$, which is significantly different than the shape of a typical Serkowski curve (Fig. 1 upper right panel). Although the difference appears to be significant, it would be challenging to identify it in practice, because multi-band polarization measurements are usually limited to a narrow range of frequencies.
@@
Figure 1: Demonstration of the effect of 3D structure in the Serkowski realtion, using analytical models for $P-\lambda$ and $\theta-\lambda$ curves. In each case, the solid black curve corresponds to the combined model and the dashed red and dotted blue lines corresponds to individual clouds. Left column: The simple case of 1 cloud (Eq. 1). Middle column: An example with 2 Clouds in the LOS, where the combined $P-\lambda$ seems to follow the Serkowski relation, but $\theta-\lambda$ not. Right column: An example with 2 Clouds in the LOS where both $P-\lambda$ and $\theta-\lambda$ profiles deviate from Serkowski expectations. The parameters for each model are: Left: $P_{max}=1\%$, $\lambda_{max}=0.6\mu m$, $\theta=30^{o}$. Middle: $P_{max_{1}}=0.7\%$, $\lambda_{max_{1}}=0.5\mu m$, $\theta_{1}=0^{o}$, $P_{max_{2}}=0.8\%$, $\lambda_{max_{2}}=0.7\mu m$, $\theta_{2}=60^{o}$. Right: $P_{max_{1}}=1.3\%$, $\lambda_{max_{1}}=0.4\mu m$, $\theta_{1}=0^{o}$, $P_{max_{2}}=1.4\%$, $\lambda_{max_{2}}=0.8\mu m$, $\theta_{2}=75^{o}$. For all clouds we have used $K=1.68\cdot\lambda_{max}$ (Wilking et al., 1980).
@@
Modest LOS variations in $\lambda_{max}$ can result in even more complications. In our assumed case of two clouds, a $40\%$ difference in $\lambda_{max}$ between the two clouds leads to nearly identical to the Serkowski curve profiles (middle panel in Fig. 1), although the cumulative signal by construction consists of two underlying profiles. In this case, erroneously assuming a single cloud along the LOS would lead to inaccurate constraints for the parameters of the Serkowski relation, and hence for the grain properties. Significant differences in the EVPAs of two clouds lying along the LOS, such as the one in the employed models which is close to $80$ degrees, can be revealed through the variation of the EVPA as a function of $\lambda$ (bottom middle and right panels in Fig. 1). However, obtaining a sufficient number of data points that would allow us to probe such variations is challenging [^2] Spectropolarimetry is one of the few robust methods for accurately probing $P$ or EVPA variations with wavelength (e.g., Bagnulo et al., 2017), but so far this method has been only applied to a very limited number of stars., and for this reason these data are sparse.
@@
These analytic calculations pose a challenge in the interpretation of $P(\lambda)$ observations: while the Serkowski relation would still hold for individual clouds, the majority of the results extracted from multi-band polarization is likely contaminated by the 3D structure of the ISM, leading to inaccurate dust-grain constraints. The upside of this problem is that it opens up new possibilities for studying the magnetized ISM in 3D.

### v3

# 3D interstellar medium structure challenges the Serkowski relation

@@ Abstract
Context. The Serkowski relation is the cornerstone of studies of starlight polarizations as a function of wavelength. Although empirical, its extensive use since its inception in 1975 to describe polarization induced by interstellar dust has elevated the relation to the status of an indisputable “law”: this is the benchmark against which models of interstellar dust grains are validated.
Aims. We revisit the effects of the 3D structure of the interstellar medium (ISM) on the wavelength dependence of interstellar polarization.
Methods. We use analytical models to show how the wavelength dependence of both the polarization fraction and direction is affected by the presence of multiple clouds along the line of sight (LOS). We account for recent developments in dust distribution modeling and we utilize an expanded archive of stellar polarization measurements to establish the effect of multiple clouds along the LOS. We highlight concrete examples of stars whose polarization profiles are severely affected by LOS variations in the dust grain and magnetic field properties, and we provide a recipe to accurately fit multiple cloud Serkowski models to such cases.
Results. We present, for the first time, compelling observational evidence that the 3D structure of the magnetized ISM often results in the violation of the Serkowski relation. We show that 3D effects impact interstellar cloud parameters derived from Serkowski fits. In particular, the dust size distribution in single-cloud sightlines may differ from analyses that ignore 3D effects, with important implications for dust modeling in the Galaxy.
Conclusions. Our results suggest that multiwavelength stellar polarization measurements offer an independent probe of the LOS variations in the magnetic field orientation, and thus constitute a potentially valuable new tool for the 3D cartography of the ISM. Finally, we caution that, unless 3D effects are explicitly accounted for, a poor fit to the Serkowski relation does not, by itself, constitute conclusive evidence that a star is intrinsically polarized.
@@ 1 Introduction
Dust is ubiquitous in the interstellar medium (ISM) of our Galaxy and plays a significant role in many astrophysical processes. The thermal (gray-body) emission of interstellar dust spans many frequencies (100 – 1000 GHz), making dust a powerful tracer of the morphology of ISM structures as well as a diagnostic of their physical properties, such as temperature (Hildebrand, 1983; André et al., 2010). At the same time, dust emission poses a challenge for cosmological experiments because it veils the cosmic microwave background signal over a wide range of frequencies (Planck Collaboration et al., 2016). Due to the multifaceted role of interstellar dust in astrophysics and cosmology, a detailed understanding of its nature is imperative.
The interaction of aspherical dust grains with the magnetic field permeating the ISM leads to their alignment, resulting in the polarization of light, which can be observed from ultraviolet to submillimeter wavelengths (Andersson et al., 2015). Short-wavelength (UV, optical, near-IR) polarization is induced by dichroic extinction of the starlight (which usually starts out unpolarized) by aligned dust grains. The same grains emit long-wavelength (far-IR and submillimeter) polarized thermal radiation (Whittet, 2022; Hensley & Draine, 2023). Dust polarization provides one of the most reliable and widely employed methods of tracing the properties of the ISM magnetic fields (e.g., Skalidis & Tassis, 2021; Skalidis et al., 2021, 2022; Pattle et al., 2023; Pelgrims et al., 2023, 2024).
Dust-induced stellar polarization follows a complex empirical relation with wavelength, known as the “Serkowski relation” (Serkowski et al., 1975). It expresses the polarization fraction, $P$, as a function of the wavelength, $\lambda$, through the formula
@@
with the use of three free parameters: $P_{max}$, the maximum polarization fraction observed at $\lambda_{max}$, and $K$, which quantifies the spread of the profile. The polarization angle (also referred to as electric vector position angle, EVPA, denoted here as $\theta$) is implicitly assumed to be constant with wavelength.
@@
Since its discovery by Serkowski et al. (1975), this relation has been extensively studied and confirmed by a number of authors (Wilking et al., 1980; Whittet et al., 1992, 2001; Bagnulo et al., 2017; Cikota et al., 2018). It is so well established that it is often dubbed the “Serkowski law.” It is even a standard practice in the literature to use it to extrapolate the polarization fraction in one particular wavelength from measurements made at another wavelength (Bijas et al., 2024; Neha et al., 2024; Ginski et al., 2024). Most importantly, the Serkowski relation with its complicated shape provides stringent constraints for dust grain models (Guillet et al., 2018; Hensley & Draine, 2021; Draine, 2024).
@@
The Serkowski relation, by construction, describes the polarization induced by a single cloud. For this reason, in the bibliography, whenever the formula is applied, it is implicitly assumed (even if not stated directly) that the observed (integrated) $P(\lambda)$ signal is induced by a single dominant polarization screen (cloud) along the LOS. Reality, however, is more complicated. Gaia has revealed the third dimension of the sky (depth, Gaia Collaboration et al., 2018), and since then it has become clear that in the majority of the LOSs there is more than one cloud with a comparable contribution in the total dust column (e.g., Green et al., 2019; Edenhofer et al., 2024). The variations in the magnetic field along the LOS are imprinted on the individual stellar polarization measurements, allowing for the tomographic mapping of the ISM magnetic field (Pelgrims et al., 2024).
@@
The impact of multiple polarizing clouds along the LOS on stellar polarization has been discussed in the literature, though not extensively. Most of the focus has been on the effect on the EVPA, with less attention given to the polarization fraction spectral profile. During the early stages of interstellar polarization (ISP) studies, when the notion of a wavelength-independent EVPA emerged from observational evidence, Treanor (1963) briefly mentioned the possibility of a nonconstant EVPA in scenarios involving multiple clouds with different magnetic field alignment and/or different grain sizes. Coyne & Gehrels (1966) attributed the observed variability in EVPA with wavelength among their targets to the potential presence of multiple clouds along the LOS. Martin (1974) was the first to construct mathematical models to describe such scenarios, involving multiple clouds with differing properties along the LOS. Some authors focused on the impact of multiple dust components in the LOS on the parameters derived from Eq. 1. For example, Codina-Landaberry & Magalhaes (1976) explored the behavior of the $K$ parameter in a medium with a changing grain alignment, while Clarke & Al-Roubaie (1984) investigated how the presence of two clouds along the LOS would influence the parameters $K$ and $\lambda_{max}$, finding no firm relationship between their models and contemporary observations.
@@
While modern researchers occasionally discuss scenarios of multiple clouds, most of the focus remains on the EVPA (e.g., McMillan & Tapia, 1977; Messinger et al., 1997; Clarke, 2010; Whittet, 2015; Bagnulo et al., 2017; Whittet, 2022). Little attention is given to the broader consequences of such scenarios, in either the $P-\lambda$ profile [^1] Patat et al. (2010) briefly hints about this in Appendix B. or the derived dust parameters.
@@
In this article, we revisit the effects of the 3D structure of the ISM on the wavelength dependence of ISP, accounting for recent developments in dust distribution modeling and utilizing an expanded archive of polarimetric measurements. We use analytical models to show how both the $P-\lambda$ and $\theta-\lambda$ profiles are expected to behave in scenarios with multiple clouds along the LOS. For the first time, we present compelling observational evidence to demonstrate these effects and disentangle the polarizing components within the LOS. Ultimately, we demonstrate that the LOS variations in dust and magnetic field properties affect the Serkowski relation in two ways: (1) they challenge the universality of the relation, because Eq. 1 can be a poor fit to the data when 3D effects are significant; and (2) even when Eq. 1 fits the data well, the best-fit parameters may not be representative of dust physics. Conversely, deviations of stellar polarization spectral profiles could potentially be utilized as a diagnostic of the 3D complexity of the magnetized ISM.
@@ 2 Theoretical expectations
In the case of multiple dust clouds along the LOS, initially unpolarized star light is affected sequentially by all of them, until it reaches the observer. In the limit of low polarization, the normalized Stokes parameters, $q=Q/I$, $u=U/I$, of the transversing light, are additive (e.g. Patat et al., 2010). Therefore, the observed $q,u$ at a given wavelength for an intrinsically unpolarized source lying behind $N_{C}$ number of clouds will be
$q_{obs}=\sum^{N_{C}}_{i=1}q_{i},$ (2)
and similarly for $u$, with $q_{i}$ being the $q$ induced by the $i-th$ dust cloud in the LOS.
In order to explore this effect, we created analytical models for the cases of one and two dust clouds along the LOS, for different parameters of the clouds. Given that the Serkowski relation is an intrinsic property of individual ISM clouds, variations in grain sizes would be manifested by variations in $\lambda_{max}$, which is proportional to the average size of aligned grains (Draine, 2024). Variations in the plane-of-sky (POS) morphology of the magnetic field are traceable through the polarization angle, $\theta$. Variations in $\lambda_{max}$ and in $\theta$ between two or more clouds along the LOS significantly affect the cumulative (observed) polarization signal. For our analytical models, we used Eq.1 to calculate $P_{i}(\lambda)$ for $0.1\mu m\leq\lambda\leq 1.2\mu m$, selecting typical values for $P_{max}$ and $\lambda_{max}$ and fixing $K=1.68\cdot\lambda_{max}$ (Wilking et al., 1980). We also assigned a $\theta_{i}$ value (constant with wavelength) representing the average POS magnetic field direction within the cloud. Thus, for each cloud, $i$, we have $q_{i}(\lambda)=P_{i}(\lambda)\cdot\cos{2\theta_{i}}$ and $u_{i}(\lambda)=P_{i}(\lambda)\cdot\sin{2\theta_{i}}$. For the combined model, we calculated $q_{obs}(\lambda)$, $u_{obs}(\lambda)$ from Eq. 2 and finally $\hat{P}(\lambda)$, $\hat{\theta}(\lambda)$ as
@@
$\hat{\theta}(\lambda)=\frac{1}{2}\cdot\tan^{-1}\left(\frac{u_{obs}(\lambda)}{q_{obs}(\lambda)}\right),$ (3b)
where we used the two-argument arctan function.
@@
In Fig. 1, we present some examples of $P-\lambda$ and $\theta-\lambda$ profiles resulting from the presence of one or two clouds in the LOS, explored by our analytical models. The simple case of a single dust cloud (Serkowski relation - Eq.1) is shown in the left panel of Fig. 1 for reference. Three cases with two clouds in the LOS are shown in the middle left, middle right, and right panels of Fig. 1. Specifically,the middle left panel demonstrates an example where both the $P-\lambda$ and $\theta-\lambda$ curves resemble the Serkowski relation, although the profiles are constructed by two components. The middle right panel demonstrates an example where the $P-\lambda$ curve resembles the Serkowski relation, yet the $\theta-\lambda$ profile deviates from constancy. The right panel demonstrates an example where both the $P-\lambda$ and the $\theta-\lambda$ profiles deviate from the Serkowski curves.
@@
Assuming a difference of a factor of two in $\lambda_{max}$ between our two hypothetical clouds yields a bimodal $P(\lambda)$, which is significantly different than the shape of a typical Serkowski curve (Fig. 1 upper right panel). Although the difference appears to be significant, it would be challenging to identify it in practice without the use of spectropolarimetry. Current polarimetric research on the ISM predominantly employs wide-filter polarimetry, which has limited resolving power. Moreover, studies with multiband polarization measurements (such as the ones used in this work - see Sect. 3.1) are usually limited to a narrow range of frequencies.
@@
Figure 1: Demonstration of the effect of 3D structure in the Serkowski relation, using analytical models for $P-\lambda$ and $\theta-\lambda$ curves. In each case, the solid black curve corresponds to the combined model and the dashed red and dotted blue lines correspond to individual clouds. Left column: Simple case of one cloud (Eq. 1). Middle left column: Example with two clouds in the LOS, in which both the combined $P-\lambda$, and $\theta-\lambda$ seem to follow the Serkowski relation. Middle right column: Example with two clouds in the LOS, in which the combined $P-\lambda$ seems to follow the Serkowski relation, but $\theta-\lambda$ not. Right column: Example with two clouds in the LOS in which both $P-\lambda$ and $\theta-\lambda$ profiles deviate from Serkowski expectations. The parameters for each model are: Left: $P_{max}=1\%$, $\lambda_{max}=0.6\mu m$, $\theta=30^{o}$. Middle left: $P_{max_{1}}=1.5\%$, $\lambda_{max_{1}}=0.55\mu m$, $\theta_{1}=30^{o}$, $P_{max_{2}}=0.7\%$, $\lambda_{max_{2}}=0.5\mu m$, $\theta_{2}=35^{o}$. Middle right: $P_{max_{1}}=0.7\%$, $\lambda_{max_{1}}=0.5\mu m$, $\theta_{1}=0^{o}$, $P_{max_{2}}=0.8\%$, $\lambda_{max_{2}}=0.7\mu m$, $\theta_{2}=60^{o}$. Right: $P_{max_{1}}=1.3\%$, $\lambda_{max_{1}}=0.4\mu m$, $\theta_{1}=0^{o}$, $P_{max_{2}}=1.4\%$, $\lambda_{max_{2}}=0.8\mu m$, $\theta_{2}=75^{o}$. For all clouds, we have used $K=1.68\cdot\lambda_{max}$ (Wilking et al., 1980)
@@
Modest LOS variations in $\lambda_{max}$ can result in even more complications. In our assumed case of two clouds, a $40\%$ difference in $\lambda_{max}$ between the two clouds leads to a profile that is nearly identical to the single cloud Serkowski curve profile (middle right panel in Fig. 1), although the cumulative signal by construction consists of two underlying profiles. In this case, erroneously assuming a single cloud along the LOS would lead to inaccurate constraints for the parameters of the Serkowski relation, and hence for the grain properties. The imprint of a second cloud in these cases can be noticed by the nonconstant behavior of $\theta$ with wavelength.
In cases in which the $\lambda_{max}$ and $\theta$ parameters are similar between clouds (middle left panel in Fig. 1), the resulting profiles of both polarization and EVPA are indistinguishable from the single cloud case within observational uncertainties, given the current accuracy of optical polarimeters. Significant differences in the EVPAs of two clouds lying along the LOS, such as the ones employed in the middle right and rightmost models of Fig. 1, can be revealed through the variation in the EVPA as a function of $\lambda$. However, obtaining a sufficient number of data points that would allow us to probe such variations is challenging [^2] Spectropolarimetry is one of the few robust methods of accurately probing $P$ or EVPA variations with wavelength (e.g., Bagnulo et al., 2017), but so far this method has only been applied to a very limited number of stars., and for this reason these data are sparse.
@@
These analytic calculations pose a challenge in the interpretation of $P(\lambda)$ observations: while the Serkowski relation would still hold for individual clouds, the majority of the results extracted from multiband polarization are likely contaminated by the 3D structure of the ISM, leading to inaccurate dust-grain constraints. The upside of this problem is that it opens up new possibilities for studying the magnetized ISM in 3D.